### v3

@@ 6.1 Basic Facts and Definitions
First note that, if there are no crossings that involve virtual edges in the embeddings of the skeletons, then the embeddings are good and their 2-clique sum yields an outer-1-planar embedding of $G$ by Section 6.2.1. If there are virtual edges that are involved in a crossing, they belong to the skeletons of P-nodes by Conditions 2 and 3. We reduce to the case of good embeddings by extending such P-nodes, that is, by modifying the SPQR-tree, its skeletons, and their embeddings as follows; refer to Figure 13. Let $e^{\prime},e^{\prime\prime}$ be two crossing virtual edges in an embedding of a skeleton of a P-node $\mu$. Let $\{u,v\}$ be the vertices of $\mathrm{skel}(\mu)$. By Condition 4, we have that the nodes $\nu^{\prime}=\mathrm{refn}(e^{\prime})$ and $\nu^{\prime\prime}=\mathrm{refn}(e^{\prime\prime})$ are S-nodes with skeletons $(u,c^{\prime}_{1},\dots,c^{\prime}_{k},v,u)$ and $(u,c_{1}^{\prime\prime},\dots,c_{\ell}^{\prime\prime},v,u)$, respectively. Without loss of generality, we may assume that the segment of $e^{\prime}$ incident to $u$ is incident to the outer face in the embedding of $\mathrm{skel}(\mu)$, and hence by Condition 1 the segment of $e^{\prime\prime}$ incident to $v$ is also incident to the outer face. By Condition 4, the edge $\{c_{k}^{\prime},v\}$ of $\mathrm{skel}(\nu^{\prime})$, which is the first edge of $e^{\prime}$ at $v$, and the edge $\{u,c_{1}^{\prime\prime}\}$ of $\mathrm{skel}(\nu^{\prime\prime})$, which is the first segment of $e^{\prime\prime}$ at $u$, are real edges. We perform the following modifications. First, we shorten the skeletons of $\nu^{\prime}$ and of $\nu^{\prime\prime}$ to $(u,c^{\prime}_{1},\dots,c^{\prime}_{k-1},c^{\prime}_{k},u)$ and $(c^{\prime\prime}_{1},c^{\prime\prime}_{2},\dots,c^{\prime\prime}_{\ell},v,c^{\prime\prime}_{1})$, respectively, and keep their embeddings planar. Second, we subdivide the virtual edge $e^{\prime}$ ($e^{\prime\prime}$) with the vertex $c^{\prime}_{k}$ ($c^{\prime\prime}_{1}$) such that $\{v,c^{\prime}_{k}\}$ ($\{u,c^{\prime\prime}_{1}\}$) is a real edge and $\{c^{\prime}_{k},u\}$ ($\{c^{\prime\prime}_{1},v\}$) is a virtual edge that is refined by the modified S-node $\nu^{\prime}$ ($\nu^{\prime\prime}$). We embed the skeletons of such nodes so that the real edges $\{v,c^{\prime}_{k}\}$ and $\{u,c^{\prime\prime}_{1}\}$ cross and the two virtual edges $\{u,c^{\prime}_{k}\}$ and $\{v,c^{\prime\prime}_{1}\}$ are incident to the outer face. Note that this essentially moves two real edges from S-node skeletons into an adjacent P-node. See the right side of Figure 13 for the modified skeletons.
@@
Observe that this construction can be carried out in linear time. Conversely, given an outer-1-planar embedding $\mathcal{E}$ of $G$. The embeddings $\mathcal{E}_{\mu}$ of all skeletons $\mathrm{skel}(\mu)$ can be obtained by checking the order of the incident edges for each vertex and the pairs of edges that cross. Moreover, it is then clear that applying the above construction to the $\mathcal{E}_{\mu}$, we reobtain $\mathcal{E}$.
The tree $\mathcal{T}^{\star}$ we constructed towards the end of the proof of Theorem 6.6 is called the extension of $\mathcal{T}$ with respect to $\{\mathcal{E}_{\mu}\}_{\mu\in V(\mathcal{T})}$. Since the modification neither adds nor removes nodes of $\mathcal{T}$, there is a bijection between the nodes of $\mathcal{T}$ and $\mathcal{T}^{*}$. Namely, each node $\mu$ of $\mathcal{T}$ contributes a unique node $\mu^{*}$ to $\mathcal{T}^{*}$, and these are all the nodes of $\mathcal{T}^{*}$. The node $\mu^{*}$ is called the extended node of $\mu$ and the good embedding of $\mathrm{skel}(\mu^{*})$ obtained with the procedure described above is called the extended embedding of $\mathrm{skel}(\mu)$. We further note that for P- and R-nodes $\mu$$\mathrm{skel}(\mu^{\star})$ only depends on the embedding $\mathcal{E}_{\mu}$ of $\mathrm{skel}(\mu)$. Thus, for P-nodes and R-nodes $\mathrm{skel}(\mu^{\star})$ can also be defined if we only have an embedding of $\mathrm{skel}(\mu)$ that satisfies 14. Condition 5 is only needed if we want to combine such skeletons into the whole extension $\mathcal{T}^{\star}$.
@@
To keep track of the existence of directed paths, we orient some edges of the skeletons of $\mathcal{T}$. Consider a skeleton $\mathrm{skel}(\mu)$ and let $e=\{u,v\}$ be an edge of $\mathrm{skel}(\mu)$. We orient $e$ from $u$ to $v$ if either it is a real edge directed from $u$ to $v$ in $G$ or if $\mathrm{expn}(e)$ contains a directed path from $u$ to $v$.
We observe some properties of the constructed orientation. First, note that a virtual edge $\{u,v\}$ is oriented in at most one direction, but it may also remain undirected in case its expansion graph contains neither a directed path from $u$ to $v$ nor a directed path from $v$ to $u$. Second, if a virtual edge $e$ with endpoints $u,v$ is directed, say from $u$ to $v$, then $u$ is the single source of $\mathrm{expn}(e)$, as otherwise $G$ would contain an additional source. Finally, if $e$ is undirected, since $G$ has a single source $s$, either $\mathrm{expn}(e)$ contains $s$ as a non-pole vertex or both $u$ and $v$ are sources of $\mathrm{expn}(e)$.
@@
Suppose that the edges of the skeletons are directed as described above, let $\mu$ be a node with an embedding $\mathcal{E}_{\mu}$ that satisfies C1–C4. Recall from the remark after the proof of Theorem 6.6 that $\mathrm{skel}(\mu^{\star})$ and its embedding $\mathcal{E}_{\mu}^{\star}$ depend only on $\mathcal{E}_{\mu}$. We call $\mathcal{E}_{\mu}$ acyclic if the planarization of the extended embedding $\mathcal{E}_{\mu}^{\star}$ is acyclic; see Figure 14 for examples. It is not surprising that indeed such embeddings are necessary for the existence of an upward outer-1-planar embedding of $G$.
Lemma 6.8 (Necessity).
Let $\mathcal{E}$ be an upward outer-1-planar embedding of $G$ without unnecessary crossings. Then the embedding of $\mathrm{skel}(\mu)$ induced by $\mathcal{E}$ is acyclic for each P- or R-node $\mu$ of the SPQR-tree $\mathcal{T}$ of $G$.
Proof 6.9.
By Theorem $\mathcal{E}$ induces embeddings $\mathcal{E}_{\mu}$ for each skeleton $\mathrm{skel}(\mu)$ of $\mathcal{T}$, which satisfy the Conditions 15. Let $\mathcal{T}^{*}$ be the extension of $\mathcal{T}$ with respect to these embeddings and recall from the proof of Theorem 6.6 that $\mathcal{E}$ can be obtained as the 2-clique sum of the embeddings $\mathcal{E}_{\mu}^{\star}$. Let $\mu$ be a P- or R-node of $\mathcal{T}$ with embedding $\mathcal{E}_{\mu}$ and let $\mu^{*}$ be the extended node in $\mathcal{T}^{*}$ with embedding $\mathcal{E}_{\mu}^{*}$. As the only crossings in $\mathcal{E}_{\mu}^{*}$ are between real edges, we can form its planarization by replacing each crossing by a dummy vertex, thereby subdividing the involved directed edges into paths of length 2. Since the upward embedding $\mathcal{E}$ in particular induces an upward embedding of its planarization, it follows that the planarization of $\mathcal{E}_{\mu}^{\star}$ does not contain a cycle that consists of directed edges, i.e., $\mathcal{E}_{\mu}$ is acyclic.
Lemma 6.10 (Sufficiency).
@@
Let $\mathcal{T}^{\star}$ be the extension of $\mathcal{T}$ with respect to the embeddings $\{\mathcal{E}_{\mu}\}_{\mu\in V(\mathcal{T})}$ of the skeletons and let $\mathcal{E}$ denote the combined embedding of $\{\mathcal{E}_{\mu}\}_{\mu\in V(\mathcal{T})}$. By Theorem 6.6, $\mathcal{E}$ is an outer-1-planar embedding of $G$. In the remainder, we show that its planarization $G^{\prime}$ is acyclic. Thus, assume for the sake of contradiction that $G^{\prime}$ contains a directed cycle $C$.
@@
We will consider the projection of $C$ to the planarizations of the skeletons of the nodes of $\mathcal{T}^{\star}$. Let $\mu^{\star}$ be a node of $\mathcal{T}^{\star}$ and consider its planarization $\mathrm{skel}^{+}(\mu^{\star})$. An edge $e$ of $\mathrm{skel}^{+}(\mu^{\star})$ belongs to the projection of $C$ if either (i) $e$ is a real edge that belongs to $C$ or (ii) $e$ is a virtual edge whose expansion graph contains a real edge that belongs to $C$. Observe that for each planarized skeleton, the projection of $C$ is either a cycle $C^{\prime}$ or a single edge. In the former case, we also say that $C$ projects to the cycle $C^{\prime}$.
@@
Claim: There exists a P- or R-node $\mu^{\star}$ of $\mathcal{T}^{\star}$ where the cycle $C$ projects to a cycle $C^{\prime}$ in the planarization $\mathrm{skel}^{+}(\mu^{\star})$ such that the source $s$ of $G$ is either a vertex of $\mathrm{skel}^{+}(\mu^{\star})$ or belongs to the expansion graph of an edge of $\mathrm{skel}^{+}(\mu^{\star})$ that does not belong to $C^{\prime}$.
@@
To prove the claim, observe that the skeletons where $C$ projects to a cycle form a subtree $\mathcal{T}_{C}$ of $\mathcal{T}^{\star}$. Similarly, the nodes whose skeletons contain $s$ as a vertex also form a subtree $\mathcal{T}_{s}$ of $\mathcal{T}^{\star}$. We choose $\mu^{\star}$ either as a common node of these two subtrees (if one exists) or as the node of $\mathcal{T}_{C}$ that is closest to the subtree $\mathcal{T}_{s}$. In the former case, we have that $s$ is a vertex of $\mathrm{skel}^{+}(\mu^{\star})$. Note that, since $s$ is a source, it is not contained in $C$ and therefore it is also not contained in the projection of $C$ to $\mathrm{skel}^{+}(\mu^{\star})$. In particular $\mu^{\star}$ cannot be an S-node. In the latter case let $e$ denote the virtual edge of $\mathrm{skel}^{+}(\mu^{\star})$ whose expansion graph contains $s$. If $e$ is contained in the projection of $C$ to $\mathrm{skel}^{+}(\mu^{\star})$, then $\mathrm{refn}(e)$ belongs to $\mathcal{T}_{C}$ and is closer to $\mathcal{T}_{s}$ than $\mu^{\star}$, which contradicts the choice of $\mu^{\star}$. Note that this in particular excludes the case that $\mu^{\star}$ is an S-node. This completes the proof of the claim.
@@
Consider the node $\mu^{\star}$ from the claim and let $\mu$ be the corresponding node of $\mathcal{T}$. The projection $C^{\prime}$ of $C$ to $\mathrm{skel}^{+}(\mu^{\star})$ may contain both real and virtual edges. If all edges of $C^{\prime}$ are directed, the cycle is already present in $\mathrm{skel}^{+}(\mu^{\star})$, i.e., the embedding $\mathcal{E}_{\mu}$ of $\mathrm{skel}(\mu)$ is not acyclic; a contradiction. Since real edges are always directed, it hence follows that $C^{\prime}$ contains an undirected virtual edge $\{u,v\}$ of $\mathrm{skel}(\mu^{\star})$. Since in expanded skeletons only real edges have crossings, the edge $\{u,v\}$ is uncrossed in the embedding of $\mathrm{skel}(\mu^{\star})$. Therefore $\mathrm{expn}(\{u,v\})$ is not crossed by edges outside of $\mathrm{expn}(\{u,v\})$. In particular the planarization of $\mathrm{expn}(\{u,v\})$ is connected to the rest of the graph only via the vertices $u$ and $v$ in the planarization $G^{\prime}$ of $G$. Since $\{u,v\}$ belongs to $C^{\prime}$, the planarization of $\mathrm{expn}(\{u,v\})$ contains a directed path $\pi$ between its two poles, say from $u$ to $v$. However, since $\{u,v\}$ is undirected and does not contain $s$, both $u$ and $v$ are sources in $\mathrm{expn}(\{u,v\})$ and hence they are sources also in its planarization. This contradicts the existence of $\pi$, and hence proves the lemma.
@@
In light of Lemmas 6.8 and 6.10, it suffices to test whether the skeletons of $\mathcal{T}$ admit embeddings that satisfy 15 and that are acyclic for P- and R-nodes. We note that 14 and acyclicity are local conditions that can be checked and all solutions can be enumerated [^1] Recall that P-node skeletons have at most five edges and each R-node skeleton is a $K_{4}$. locally and independently for each skeleton. On the other hand 5 is a global property, which states that for each real edge $e$ that connects two P-nodes in a series, only one of them may put a crossing on $e$.
@@
For each node $\mu$ of $\mathcal{T}$, let $\mathcal{F}_{\mu}$ be a subset of the embeddings of $\mathrm{skel}(\mu)$ that satisfy 14. We call $\mathcal{F}_{\mu}$ the feasible embeddings of $\mathrm{skel}(\mu)$. We are interested in whether we can choose for each node $\mu$ a feasible embedding $\mathcal{E}_{\mu}\in\mathcal{F}_{\mu}$ such that together they satisfy 5. We call such a choice of embeddings consistent.
@@
To find a consistent choice of embeddings, we construct an auxiliary graph $H$ whose vertex set is $\bigcup_{\mu\in\mathcal{T}}\mathcal{F}_{\mu}$. We turn each set $\mathcal{F}_{\mu}$ into a clique and we connect two embedding choices of different P-nodes if choosing both of them simultaneously violates 5. It is then clear that $H$ contains an independent set whose size equals the number of nodes of $\mathcal{T}$ if and only if there exists a consistent choice of embeddings. In the following we show that this can be decided in linear time by proving that $H$ has bounded treewidth [15].
Lemma 6.12.
@@
We now use this to bound the treewidth of $H$. For each S-node $\lambda$ of $\mathcal{T}$ the skeleton $\mathrm{skel}(\lambda)$ defines a circular ordering of its virtual edges. Note that embeddings of two different P-nodes $\mu$ and $\mu^{\prime}$ can be connected by an edge only if they have a common S-node neighbor $\lambda$ such that their corresponding virtual edges in $\mathrm{skel}(\lambda)$ are consecutive in $\mathrm{skel}(\lambda)$. Therefore the auxiliary graph $H$ is a subgraph of the graph obtained from the SPQR-tree $\mathcal{T}$ by (i) forming the closure $T^{\prime}$ of $\mathcal{T}$ with respect to the (independent set of) S-nodes, where we connect the neighbors of the S-node in the order in which the corresponding virtual edges appear in the skeleton of the S-node and (ii) taking the $c$-clique-expansion of $T^{\prime}$ where $c$ is an upper bound on the number of embeddings satisfying C1–C4 for any skeleton. Auer et al. [4] show that $c\leq 12$. Hence the treewidth of $H$ is at most $36$.
Theorem 6.14.
@@
Proof 6.15.
By Lemma 6.4 we may assume that $G$ is biconnected. We first compute the SPQR-tree $\mathcal{T}$ of $G$ in linear time [30]. Next, we check that the skeleton of each R-node of $\mathcal{T}$ is a $K_{4}$, and that the skeleton of each P-node of $\mathcal{T}$ has at most four virtual edges. If this fails, we can reject the instance as it does not have an outer-1-planar embedding [4]. Next, we compute for each node $\mu$ of $\mathcal{T}$ the set $\mathcal{F}_{\mu}$ that contains all acyclic embeddings of $\mathrm{skel}(\mu)$ that satisfy conditions 14. Since skeletons of S-nodes have a unique planar embedding (2) and the skeletons of P- and R-nodes have bounded size, this can be done in total linear time. It then remains to consistently choose these embeddings so that also Condition 5 is satisfied. To this end, we construct the auxiliary graph $H$ and compute a maximum independent set, which takes linear time as well [15]. If the size of the maximum independent set is smaller than the number of nodes of $\mathcal{T}$, there is no consistent choice of embeddings and we reject the instance. Otherwise, by Theorem 6.6 this choice defines an outer-1-planar embedding $\mathcal{E}$. By Lemma 6.10 the embedding $\mathcal{E}$ has an acyclic planarization and is hence upward outer-1-planar by Theorem 6.2. We note that, in the positive case, the embedding $\mathcal{E}$ can also be constructed in linear time.

### v4

@@ 6.2 Proof of Theorem 7
First note that, if there are no crossings that involve virtual edges in the embeddings of the skeletons, then the embeddings are good and their 2-clique sum yields an outer-1-planar embedding of $G$ by 1. If there are virtual edges that are involved in a crossing, they belong to the skeletons of P-nodes by Conditions 2 and 3. We reduce to the case of good embeddings by extending such P-nodes, that is, by modifying the SPQR-tree, its skeletons, and their embeddings as follows; refer to Fig. 13. Let $e^{\prime},e^{\prime\prime}$ be two crossing virtual edges in an embedding of a skeleton of a P-node $\mu$. Let $\{u,v\}$ be the vertices of $\operatorname{skel}(\mu)$. By Condition 4, we have that the nodes $\nu^{\prime}=\operatorname{refn}(e^{\prime})$ and $\nu^{\prime\prime}=\operatorname{refn}(e^{\prime\prime})$ are S-nodes with skeletons $(u,c^{\prime}_{1},\dots,c^{\prime}_{k},v,u)$ and $(u,c_{1}^{\prime\prime},\dots,c_{\ell}^{\prime\prime},v,u)$, respectively. Without loss of generality, we may assume that the segment of $e^{\prime}$ incident to $u$ is incident to the outer face in the embedding of $\operatorname{skel}(\mu)$, and hence by Condition 1 the segment of $e^{\prime\prime}$ incident to $v$ is also incident to the outer face. By Condition 4, the edge $\{c_{k}^{\prime},v\}$ of $\operatorname{skel}(\nu^{\prime})$, which is the first segment of $e^{\prime}$ at $v$, and the edge $\{u,c_{1}^{\prime\prime}\}$ of $\operatorname{skel}(\nu^{\prime\prime})$, which is the first segment of $e^{\prime\prime}$ at $u$, are real edges. We perform the following modifications. First, we shorten the skeletons of $\nu^{\prime}$ and of $\nu^{\prime\prime}$ to $(u,c^{\prime}_{1},\dots,c^{\prime}_{k-1},c^{\prime}_{k},u)$ and $(c^{\prime\prime}_{1},c^{\prime\prime}_{2},\dots,c^{\prime\prime}_{\ell},v,c^{\prime\prime}_{1})$, respectively, and keep their embeddings planar. Second, we subdivide the virtual edge $e^{\prime}$ ($e^{\prime\prime}$) with the vertex $c^{\prime}_{k}$ ($c^{\prime\prime}_{1}$) such that $\{v,c^{\prime}_{k}\}$ ($\{u,c^{\prime\prime}_{1}\}$) is a real edge and $\{c^{\prime}_{k},u\}$ ($\{c^{\prime\prime}_{1},v\}$) is a virtual edge that is refined by the modified S-node $\nu^{\prime}$ ($\nu^{\prime\prime}$). We embed the skeletons of such nodes so that the real edges $\{v,c^{\prime}_{k}\}$ and $\{u,c^{\prime\prime}_{1}\}$ cross and the two virtual edges $\{u,c^{\prime}_{k}\}$ and $\{v,c^{\prime\prime}_{1}\}$ are incident to the outer face. Note that this essentially moves two real edges from S-node skeletons into an adjacent P-node. See the right side of Fig. 13 for the modified skeletons.
@@
Observe that this construction can be carried out in linear time. Conversely, given an outer-1-planar embedding $\mathcal{E}$ of $G$. The embeddings $\mathcal{E}_{\mu}$ of all skeletons $\operatorname{skel}(\mu)$ can be obtained by checking the order of the incident edges for each vertex and the pairs of edges that cross. Moreover, it is then clear that applying the above construction to the $\mathcal{E}_{\mu}$, we reobtain $\mathcal{E}$. ∎
The tree $\mathcal{T}^{\star}$ we constructed towards the end of the proof of Theorem 9 is called the extension of $\mathcal{T}$ with respect to $\{\mathcal{E}_{\mu}\}_{\mu\in V(\mathcal{T})}$. Since the modification neither adds nor removes nodes of $\mathcal{T}$, there is a bijection between the nodes of $\mathcal{T}$ and $\mathcal{T}^{*}$. Namely, each node $\mu$ of $\mathcal{T}$ contributes a unique node $\mu^{*}$ to $\mathcal{T}^{*}$, and these are all the nodes of $\mathcal{T}^{*}$. The node $\mu^{*}$ is called the extended node of $\mu$ and the good embedding of $\operatorname{skel}(\mu^{*})$ obtained with the procedure described above is called the extended embedding of $\operatorname{skel}(\mu)$. We further note that for P- and R-nodes $\mu$$\operatorname{skel}(\mu^{\star})$ only depends on the embedding $\mathcal{E}_{\mu}$ of $\operatorname{skel}(\mu)$. Thus, for P-nodes and R-nodes $\operatorname{skel}(\mu^{\star})$ can also be defined if we only have an embedding of $\operatorname{skel}(\mu)$ that satisfies 14. Condition 5 is only needed if we want to combine such skeletons into the whole extension $\mathcal{T}^{\star}$.
@@
To keep track of the existence of directed paths, we orient some edges of the skeletons of $\mathcal{T}$. Consider a skeleton $\operatorname{skel}(\mu)$ and let $e=\{u,v\}$ be an edge of $\operatorname{skel}(\mu)$. We orient $e$ from $u$ to $v$ if either it is a real edge directed from $u$ to $v$ in $G$ or if $\operatorname{expn}(e)$ contains a directed path from $u$ to $v$.
We observe some properties of the constructed orientation. First, note that a virtual edge $\{u,v\}$ is oriented in at most one direction, but it may also remain undirected in case its expansion graph contains neither a directed path from $u$ to $v$ nor a directed path from $v$ to $u$. Second, if a virtual edge $e$ with endpoints $u,v$ is directed, say from $u$ to $v$, then $u$ is the single source of $\operatorname{expn}(e)$, as otherwise $G$ would contain an additional source. Finally, if $e$ is undirected, since $G$ has a single source $s$, either $\operatorname{expn}(e)$ contains $s$ as a non-pole vertex or both $u$ and $v$ are sources of $\operatorname{expn}(e)$.
@@
Suppose that the edges of the skeletons are directed as described above, let $\mu$ be a node with an embedding $\mathcal{E}_{\mu}$ that satisfies C1–C4. Recall from the remark after the proof of Theorem 9 that $\operatorname{skel}(\mu^{\star})$ and its embedding $\mathcal{E}_{\mu}^{\star}$ depend only on $\mathcal{E}_{\mu}$. We call $\mathcal{E}_{\mu}$ acyclic if the planarization of the extended embedding $\mathcal{E}_{\mu}^{\star}$ is acyclic; see Fig. 14 for examples. It is not surprising that indeed such embeddings are necessary for the existence of an upward outer-1-planar embedding of $G$.
Lemma 3 (Necessity).
Let $\mathcal{E}$ be an upward outer-1-planar embedding of $G$ without unnecessary crossings. Then the embedding of $\operatorname{skel}(\mu)$ induced by $\mathcal{E}$ is acyclic for each P- or R-node $\mu$ of the SPQR-tree $\mathcal{T}$ of $G$.
Proof.
By Theorem 9, $\mathcal{E}$ induces embeddings $\mathcal{E}_{\mu}$ for each skeleton $\operatorname{skel}(\mu)$ of $\mathcal{T}$, which satisfy the Conditions 15. Let $\mathcal{T}^{*}$ be the extension of $\mathcal{T}$ with respect to these embeddings and recall from the proof of Theorem 9 that $\mathcal{E}$ can be obtained as the 2-clique sum of the embeddings $\mathcal{E}_{\mu}^{\star}$. Let $\mu$ be a P- or R-node of $\mathcal{T}$ with embedding $\mathcal{E}_{\mu}$ and let $\mu^{*}$ be the extended node in $\mathcal{T}^{*}$ with embedding $\mathcal{E}_{\mu}^{*}$. As the only crossings in $\mathcal{E}_{\mu}^{*}$ are between real edges, we can form its planarization by replacing each crossing by a dummy vertex, thereby subdividing the involved directed edges into paths of length 2. Since the upward embedding $\mathcal{E}$ in particular induces an upward embedding of its planarization, it follows that the planarization of $\mathcal{E}_{\mu}^{\star}$ does not contain a cycle that consists of directed edges, i.e., $\mathcal{E}_{\mu}$ is acyclic. ∎
Lemma 4 (Sufficiency).
@@
Let $\mathcal{T}^{\star}$ be the extension of $\mathcal{T}$ with respect to the embeddings $\{\mathcal{E}_{\mu}\}_{\mu\in V(\mathcal{T})}$ of the skeletons and let $\mathcal{E}$ denote the combined embedding of $\{\mathcal{E}_{\mu}\}_{\mu\in V(\mathcal{T})}$. By Theorem 9, $\mathcal{E}$ is an outer-1-planar embedding of $G$. In the remainder, we show that its planarization $G^{\prime}$ is acyclic. Thus, assume for the sake of contradiction that $G^{\prime}$ contains a directed cycle $C$.
@@
We will consider the projection of $C$ to the planarizations of the skeletons of the nodes of $\mathcal{T}^{\star}$. Let $\mu^{\star}$ be a node of $\mathcal{T}^{\star}$ and consider its planarization $\operatorname{skel}^{+}(\mu^{\star})$. An edge $e$ of $\operatorname{skel}^{+}(\mu^{\star})$ belongs to the projection of $C$ if either (i) $e$ is a real edge that belongs to $C$ or (ii) $e$ is a virtual edge whose expansion graph contains a real edge that belongs to $C$. Observe that for each planarized skeleton, the projection of $C$ is either a cycle $C^{\prime}$ or a single edge. In the former case, we also say that $C$ projects to the cycle $C^{\prime}$.
@@
Claim: There exists a P- or R-node $\mu^{\star}$ of $\mathcal{T}^{\star}$ where the cycle $C$ projects to a cycle $C^{\prime}$ in the planarization $\operatorname{skel}^{+}(\mu^{\star})$ such that the source $s$ of $G$ is either a vertex of $\operatorname{skel}^{+}(\mu^{\star})$ or belongs to the expansion graph of an edge of $\operatorname{skel}^{+}(\mu^{\star})$ that does not belong to $C^{\prime}$.
@@
To prove the claim, observe that the skeletons where $C$ projects to a cycle form a subtree $\mathcal{T}_{C}$ of $\mathcal{T}^{\star}$. Similarly, the nodes whose skeletons contain $s$ as a vertex also form a subtree $\mathcal{T}_{s}$ of $\mathcal{T}^{\star}$. We choose $\mu^{\star}$ either as a common node of these two subtrees (if one exists) or as the node of $\mathcal{T}_{C}$ that is closest to the subtree $\mathcal{T}_{s}$. In the former case, we have that $s$ is a vertex of $\operatorname{skel}^{+}(\mu^{\star})$. Note that, since $s$ is a source, it is not contained in $C$ and therefore it is also not contained in the projection of $C$ to $\operatorname{skel}^{+}(\mu^{\star})$. In particular $\mu^{\star}$ cannot be an S-node. In the latter case let $e$ denote the virtual edge of $\operatorname{skel}^{+}(\mu^{\star})$ whose expansion graph contains $s$. If $e$ is contained in the projection of $C$ to $\operatorname{skel}^{+}(\mu^{\star})$, then $\operatorname{refn}(e)$ belongs to $\mathcal{T}_{C}$ and is closer to $\mathcal{T}_{s}$ than $\mu^{\star}$, which contradicts the choice of $\mu^{\star}$. Note that this in particular excludes the case that $\mu^{\star}$ is an S-node. This completes the proof of the claim.
@@
Consider the node $\mu^{\star}$ from the claim and let $\mu$ be the corresponding node of $\mathcal{T}$. The projection $C^{\prime}$ of $C$ to $\operatorname{skel}^{+}(\mu^{\star})$ may contain both real and virtual edges. If all edges of $C^{\prime}$ are directed, the cycle is already present in $\operatorname{skel}^{+}(\mu^{\star})$, i.e., the embedding $\mathcal{E}_{\mu}$ of $\operatorname{skel}(\mu)$ is not acyclic; a contradiction. Since real edges are always directed, it hence follows that $C^{\prime}$ contains an undirected virtual edge $\{u,v\}$ of $\operatorname{skel}(\mu^{\star})$. Since in expanded skeletons only real edges have crossings, the edge $\{u,v\}$ is uncrossed in the embedding of $\operatorname{skel}(\mu^{\star})$. Therefore $\operatorname{expn}(\{u,v\})$ is not crossed by edges outside of $\operatorname{expn}(\{u,v\})$. In particular the planarization of $\operatorname{expn}(\{u,v\})$ is connected to the rest of the graph only via the vertices $u$ and $v$ in the planarization $G^{\prime}$ of $G$. Since $\{u,v\}$ belongs to $C^{\prime}$, the planarization of $\operatorname{expn}(\{u,v\})$ contains a directed path $\pi$ between its two poles, say from $u$ to $v$. However, since $\{u,v\}$ is undirected and does not contain $s$, both $u$ and $v$ are sources in $\operatorname{expn}(\{u,v\})$ and hence they are sources also in its planarization. This contradicts the existence of $\pi$, and hence proves the lemma. ∎
@@
In light of Lemmas 3 and 4, it suffices to test whether the skeletons of $\mathcal{T}$ admit embeddings that satisfy 15 and that are acyclic for P- and R-nodes. We note that 14 and acyclicity are local conditions that can be checked and all solutions can be enumerated [^1] Recall that P-node skeletons have at most five edges and each R-node skeleton is a $K_{4}$. locally and independently for each skeleton. On the other hand 5 is a global property, which states that for each real edge $e$ that connects two P-nodes in a series, only one of them may put a crossing on $e$.
@@
For each node $\mu$ of $\mathcal{T}$, let $\mathcal{F}_{\mu}$ be a subset of the embeddings of $\operatorname{skel}(\mu)$ that satisfy 14. We call $\mathcal{F}_{\mu}$ the feasible embeddings of $\operatorname{skel}(\mu)$. We are interested in whether we can choose for each node $\mu$ a feasible embedding $\mathcal{E}_{\mu}\in\mathcal{F}_{\mu}$ such that together they satisfy 5. We call such a choice of embeddings consistent.
@@
To find a consistent choice of embeddings, we construct an auxiliary graph $H$. The vertex set of $H$ is $\bigcup_{\mu\in\mathcal{T}}\mathcal{F}_{\mu}$. We turn each set $\mathcal{F}_{\mu}$ into a clique and we connect two embedding choices of different P-nodes if choosing both of them simultaneously violates 5. It is then clear that $H$ contains an independent set whose size equals the number of nodes of $\mathcal{T}$ if and only if there exists a consistent choice of embeddings. In the following, we show that this can be decided in linear time by proving that $H$ has bounded treewidth (Cygan, Fomin, Kowalik, Lokshtanov, Marx, Pilipczuk, Pilipczuk, and Saurabh, 2015).
Lemma 5.
@@
We now use this to bound the treewidth of $H$. For each S-node $\lambda$ of $\mathcal{T}$ the skeleton $\operatorname{skel}(\lambda)$ defines a circular ordering of its virtual edges. Note that embeddings of two different P-nodes $\mu$ and $\mu^{\prime}$ can be connected by an edge only if they have a common S-node neighbor $\lambda$ such that their corresponding virtual edges in $\operatorname{skel}(\lambda)$ are consecutive in $\operatorname{skel}(\lambda)$. Therefore the auxiliary graph $H$ is a subgraph of the graph obtained from the SPQR-tree $\mathcal{T}$ by (i) forming the closure $T^{\prime}$ of $\mathcal{T}$ with respect to the (independent set of) S-nodes, where we connect the neighbors of the S-node in the order in which the corresponding virtual edges appear in the skeleton of the S-node and (ii) taking the $c$-clique-expansion of $T^{\prime}$ where $c$ is an upper bound on the number of embeddings satisfying C1–C4 for any skeleton. Auer et al. (2016) show that $c\leq 12$. Hence the treewidth of $H$ is at most $36$. ∎
Theorem 10.
@@
Proof.
By Lemma 2 we may assume that $G$ is biconnected. We first compute the SPQR-tree $\mathcal{T}$ of $G$ in linear time (Gutwenger and Mutzel, 2000). Next, we check that the skeleton of each R-node of $\mathcal{T}$ is a $K_{4}$, and that the skeleton of each P-node of $\mathcal{T}$ has at most four virtual edges. If this fails, we can reject the instance as it does not have an outer-1-planar embedding (Auer et al., 2016). Next, we compute for each node $\mu$ of $\mathcal{T}$ the set $\mathcal{F}_{\mu}$ that contains all acyclic embeddings of $\operatorname{skel}(\mu)$ that satisfy conditions 14. Since skeletons of S-nodes have a unique planar embedding (2) and the skeletons of P- and R-nodes have bounded size, this can be done in total linear time. It then remains to consistently choose these embeddings so that also Condition 5 is satisfied. To this end, we construct the auxiliary graph $H$ and compute a maximum independent set, which takes linear time as well (Cygan et al., 2015). If the size of the maximum independent set is smaller than the number of nodes of $\mathcal{T}$, there is no consistent choice of embeddings and we reject the instance. Otherwise, by Theorem 9 this choice defines an outer-1-planar embedding $\mathcal{E}$. By Lemma 4 the embedding $\mathcal{E}$ has an acyclic planarization and is hence upward outer-1-planar by Theorem 8. We note that, in the positive case, the embedding $\mathcal{E}$ can also be constructed in linear time. ∎